# Towards Autonomous Reinforcement Learning for Real-World Robotic Manipulation with Large Language Models

### v2

@@ I Introduction
The recent development of Large Language Models (LLMs), and Visual Language Models (VLMs) has drastically impacted academic and industrial research, due to their flexibility to many human-like tasks [1]. LLMs have found numerous applications in robotics [2], including high-level semantic planning [3, 4, 5, 6] and low-level locomotion and manipulation skills [7, 8, 9].
The impact of LLMs and VLMs on robotic applications has opened new possibilities for how machines comprehend and interact with their environments. While these models demonstrate impressive high-level reasoning and language-based planning capabilities, the field of Reinforcement Learning (RL) [10] complements them by focusing on direct interaction with physical tasks. Due to its versatility and adaptability, RL has achieved astounding results in robotics [11] and other fields [12], enabling agents to autonomously learn and optimize complex behaviors.
The main goal of RL algorithms is to generate and optimize autonomous agents to complete tasks, by autonomously interacting with the available environment. The agent receives feedback that evaluates its performance in the form of a numerical reward, and is trained to produce actions that maximize the expected rewards. RL approaches greatly reduce the effort required to manually define complicated tasks in environments with complex dynamics, often resulting in better performances than hand-coded approaches. Nonetheless, careful design and engineering are required to define rewards to effectively and efficiently train RL agents. Many real-world tasks can be encoded by sparse rewards, which are in general easy to define but result challenging to optimize with RL. Instead, in robotics and other fields dense rewards are highly adopted in robotics to guide the exploration of autonomous agents [13] with the drawback of a more elaborate design. Indeed, dense reward functions can be difficult to tune, causing numerical instability, and misalignment w.r.t. the task objectives. Thus, automated or guided reward generation methods are interesting for both academic research and practical reasons.
@@
Several recent publications have tried to handle the design of dense reward functions in an automated fashion by using state-of-the-art coding LLMs, such as GPT-4 [14]. Their remarkable skills in code writing using natural language context allow non-expert users to define goals for RL agents as simple human-readable text. Such approaches reached remarkable performances in RL reward design in dexterous robotics manipulation [15, 16, 17, 18, 19], locomotion [20, 21], navigation [22], and autonomous driving [23]. Other applications of LLMs and VLMs for RL in robotics include approaches to guide the exploration process of RL agents [24], imitation [25, 26], online fine-tuning [27], and Preference-based RL (PbRL) [28]. Remarkably, [29] employs a VLM as the backbone of a Deep policy, and RL training is used to fine-tune the policy performance for various tasks in simulated environments. Current solutions for robotics manipulation, including [15, 7, 17, 18, 19], demonstrate few examples of applications in real-world settings, most requiring multiple training stages using feedback from human experts.
In this work, we propose Autonomous Reinforcement learning for Complex Human-Informed Environments (ARCHIE). ARCHIE is a practical automatic RL pipeline for training autonomous agents for robotics manipulation tasks, in an unsupervised manner. ARCHIE employs GPT-4 —a popular pre-trained LLM— for reward generation from human prompts. We leverage natural language descriptions to generate reward functions via GPT-4, which are then used to train an RL agent in a simulated environment. Our approach introduces a formalization of the reward function that constrains the language model’s code generation, enhancing the feasibility of task learning at the first attempt. Unlike previous methods, we also utilize the language model to define the success criteria for each task, further automating the learning pipeline. Moreover, by properly formalizing the reward functions in shaping and terminal terms, we avoid the need for reward reflection [15] and multiple stages of training in RL [18]. This results in a streamlined, one-shot process translating the user’s text descriptions into deployable skills. ARCHIE relies on the reasonable assumption that the manipulator’s kinematic and dynamic models are available and that RL agents are trained in a simulated environment. We demonstrate the effectiveness of the reward generation in a simplified simulated environment and with a selection of real-world single-arm and bi-manual manipulation tasks on an ABB YuMi collaborative robot, shown in Figure 5. Policies trained in simulation are demonstrated on the real robot in laboratory settings111videos available at https://youtu.be/kaTxZ4oqD6Y.
@@ II-A LLM-assisted RL
In [24], the authors propose the use of a VLM and an LLM to guide policy learning in the context of robotics manipulation. The VLM is used as an open-vocabulary object detector to get the observation space of the RL policy, while the action space is defined w.r.t. the reference frame of the objects. The LLM is used to program both a high-level and a low-level policy to generate affordance maps [3] according to the visual input. The LLM policies are used during the training of a Soft Actor-Critic (SAC) [30] agent with an $\varepsilon$-greedy exploration approach.
In [25, 26], the authors use the LLMs to define expert agents to guide the RL training with imitation. The approach in [25] relies on the availability of pick/place primitives that make the RL agent’s action space discrete. The LLM policies generate expert demonstration data that bootstrap the training of an SDQfD [31] agent. In [26], a variation of Twin Delayed Deep Deterministic Policy Gradient (TD3) [32] is proposed, to accommodate an imitation loss in the actor’s loss. The imitation data is collected during the training process with a robot controller programmed with an LLM.
In [27] the researchers used Cal-QL [33], a two-step policy training composed of (i) an offline stage where a policy is pre-trained to imitate various demonstrations, (ii) an online stage where the policy is fine-tuned to solve specific tasks while using a VLM to generate rewards directly from visual observations.
In [28] authors incorporate PEBBLE [34] with natural language user feedback on the iteratively learned robot policies during the PbRL training. Preferences of multiple LLMs are collected to compute a scoring of the performance of agents.
@@ II-B LLM for reward generation
Most notably, Eureka [15] is one of the most recent contributions in reward generation for RL. It employs GPT-4 to generate dense reward functions in the form of Python code by using the environment’s source code and a natural-language task description as context. Then, it iterates the training and reward reflection by querying the LLM with feedback from the previous episode to improve the performance of the agents. Such a strategy has proven to be effective in simulation, where it can accomplish various robotic tasks that cannot be solved in a sparse reward setting. Remarkably, the reward functions generated by GPT-4 prove to be more robust than rewards generated by humans. Recently, Eureka was applied in a real-world robotic locomotion task [35], but to the best of our knowledge, not to robotic manipulation.
Similarly, in [18] and [19], the authors use an LLM to design reward functions based on natural language descriptions of the task and environment. The model’s initial reward function design is then used to train an RL agent and the results are fed back to the LLM for self-refinement. This iterative process continues until the reward function meets the desired performance criteria. Notably, [19] presented results in real-world dexterous manipulation tasks.
In [16], authors use Offline RL [36] to train language-conditioned robot behaviors from large annotated offline datasets that outperform language-conditioned imitation learning techniques. The method exhibits zero-shot generalization to unseen natural language commands in real-world settings.
The work in [17] leverages LLMs and VLMs to extract knowledge about object semantics and common-sense task relevance to autonomously and continuously generate new skills through a self-guided learning cycle. In this cycle, the VLM-powered agent autonomously decides the learning task and prepares a natural language description followed by a cost or reward function. Remarkably, it also generates an appropriate simulated environment using a dictionary of available objects. Tasks are solved with RL, Model Predictive Control (MPC) [37] or other Optimal Control strategies. With this approach, the agent builds a plethora of skills for a variety of self-proposed tasks.
@@ III Notations
RL algorithms aim at optimizing the policy $\pi$ to maximizes the expected discounted sum of rewards $R_{0}=\mathbb{E_{\pi_{\phi}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$ of a specific MDP. To do so, two main tools are defined, the value function $V$ and the action-value function $Q$:
$V^{\pi}(s_{t})$ $=\mathbb{E_{\pi}}\left[R_{t}|S_{t}=s_{t}\right]$ (1)
@@ IV-B Initialization: Natural Language to Rewards
Given a task for the robot, the LLM is used to write a new Python class that extends the base class by defining the dense reward function to guide the agent’s exploration towards the task goal. Unlike previous publications, we require the LLM to implement the termination conditions for the defined task including a success condition that evaluates whether the task is solved and a failure condition that assesses whether the agent has failed. This is done to reduce the ambiguity in task definitions in natural language, as the geometrical context is very complex to infer from the environment’s code. Moreover, the failure condition reduces the amount of exploration required by the agents, since failing states are associated with low rewards and are not interesting to explore. In this way, the agent is encouraged to avoid failure instead.
@@
  • Introduction, in which the overall LLM’s task is described, including the code-writing requirements;
@@
  • Context, including the code and documentation of the base Python class. The documentation is an important part of the context as it explains how to use the APIs to register contacts and poses of the objects useful for defining the reward components.
  • General indications on how reward functions for manipulation tasks are written, including some examples.
The prompt design is reported in the appendix.
@@ IV-C Autonomous Skill Learning: Dense reward formalization
Let $r^{k}(s_{t},a_{t})$, $k=1,\dots,K$ be the $k$-th reward component of our dense reward function shaping, received after applying action $a_{t}$ in state $s_{t}$. These terms are generated by the LLM using the prompt engineering defined in Section IV-B. Arbitrarily, the reward components $r^{k_{+}}(s_{t},a_{t})$, $k_{+}\in\boldsymbol{K_{+}}\subseteq\{1,..,K\}$ are positive and called bonuses. Due to the available context, the shaping terms are, generally, coherent with the task requirements. Nonetheless, there is no guarantee that the different generated terms are numerically coherent with each other, as discussed in other works. As an example, for a pushing task, where the agent is required to move an object from a starting position to a target position, one possible reward function that LLMs are capable of generating is:
where $d$ is the distance between the object and the target position, while $b\in\mathbb{N^{+}}$ is a bonus for touching the object. Indeed, the bonus $b$ can guide exploration toward completing the desired task; however, even this simple reward requires numerical tuning. Namely, if $b$ is very big, compared to $d$, then the policy learning might learn sub-optimally since the reward landscape is flat, w.r.t. the range of the reward. We consider a numerical example of the reward in eq. 4. Namely, we define a simple 2D environment: the agent is a point on a plane, its action space is its cartesian velocity on the plane, and the target is the plane’s origin. The bonus $b$ is a constant term, i.e. the object is attached to the agent. We train a SAC agent under two rewards, the first with $b=10$, the second with $b=1$. We show the landscapes of the two rewards and the training success rate curves in Figure 2. With $b=10$, the reward landscape does not discriminate well points, i.e. rewards in the area are almost identical. Instead, with $b=1$, the goal is clearly defined, which is reflected in the performance of policy learning. Therefore, even in the presence of the correct reward terms, if the weights are not tuned correctly, the performance of the policy is severely affected.
Figure 2: Top: Rewards landscapes for two example rewards, defined in eq. 4, with constant bonus $b=10$ and $b=1$. Bottom: success rates of the SAC agents trained with the two rewards, statistics show median and 25-75 quantiles of 10 random seeds.
Indeed, authors in [15, 18] propose iterative processes where the LLM evaluates the results of the trained agents and adjusts the weights of the different addends. In an effort to avoid this iterative process, we introduce a clearly defined terminal reward term $R_{F}(s_{t},a_{t})$ and two binary classifiers $\Phi(s_{t})$ and $\Gamma(s_{t},t)$, defined as:
$\Phi(s_{t})$ =$=\begin{cases}1\text{ If in }s_{t}\text{ the task is solved}\\ 0\text{ otherwise}\end{cases}$ (5)
$\Gamma(s_{t},t)$ =$=\begin{cases}1\text{ If in }s_{t}\text{ the task is failed or }t=T-1\\ 0\text{ otherwise}\end{cases}$ (6)
Both are implemented by the LLM in the same Python class.
The reward function is defined as
$r(s_{t},a_{t},s_{t+1})=\sum_{k=1}^{K}r^{k}(s_{t},a_{t})+R_{F}(s_{t},a_{t})\cdot\Phi(s_{t+1})$ (7)
@@
To ensure that the terminal reward is sufficiently high in any terminal state $s_{t^{\prime}}$, the sufficient condition is
$r^{k_{+}}(s_{t^{\prime}},a_{t^{\prime}})\geq r^{k_{+}}(s_{t},a_{t})\quad\forall t=0,\dots,t^{\prime},\quad\forall{k_{+}}\in\boldsymbol{K_{+}}$ (9)
@@
The statement is automatically true if $\sum_{k_{+}\in\boldsymbol{K_{+}}}r^{k}(s_{t^{\prime}},a_{t^{\prime}})\leq 1$, otherwise it is sufficient to apply eq. 9 to the definition in eq. 8:
@@
$\sum_{t=0}^{t^{\prime}}\sum_{k_{+}\in\boldsymbol{K_{+}}}r^{k_{+}}(s_{t},a_{t})$ $>\sum_{t=0}^{t^{\prime}}\sum_{k=1}^{K}r^{k}(s_{t},a_{t})=\sum_{t=0}^{t^{\prime}}r(s_{t},a_{t})$ (11)
$\Box$
In this way, we guarantee that the terminal reward for completing the task is always at least one order of magnitude greater than the cumulative rewards the agent can collect during the trajectory before termination. For $\gamma<1$, this encourages the agent to move towards the terminal states rather than stagnating in high-rewarding non-terminal states, which is likely to happen if the shaping components are not balanced.
@@ IV-D Autonomous Skill Learning: Policy Learning
The pseudocode of the policy learning algorithm is reported in Algorithm 1, it follows the notations used in [30, 32] and related research. In the simulated environment, the agent’s state is defined as a vector containing the pose and linear and angular velocities of the robot’s end-effectors, the position and velocity of the grippers’ joint, and the pose of the objects in the scene. The agent’s action is defined as the reference linear and angular velocities of the end-effectors and the reference velocity of the grippers’ joint. This definition of state and action spaces allows the agent to directly correlate actions to change in the state and reward. For single-arm tasks, the state and action spaces are limited to drive the left arm, while the right arm is fixed to a constant joint configuration.

### v4

@@ I Introduction
The recent development of Large Language Models (LLMs), and Visual Language Models (VLMs) has drastically impacted academic and industrial research, due to their flexibility to many human-like tasks [1]. LLMs have found numerous applications in robotics [2], including high-level semantic planning [3, 4, 5, 6] and low-level locomotion and manipulation skills [7, 8, 9]. The impact of LLMs and VLMs on robotic applications has opened new possibilities for how machines comprehend and interact with their environments. While these models demonstrate impressive high-level reasoning and language-based planning capabilities, the field of Reinforcement Learning (RL) [10] complements them by focusing on direct interaction with physical tasks. Due to its versatility and adaptability, RL has achieved astounding results in robotics [11] and other fields [12], enabling agents to autonomously learn and optimize complex behaviors.
The main goal of RL algorithms is to generate and optimize autonomous agents to complete tasks, by autonomously interacting with the available environment. The agent receives feedback that evaluates its performance in the form of a numerical reward, and is trained to produce actions that maximize the expected rewards. RL approaches greatly reduce the effort required to manually define complicated tasks in environments with complex dynamics, often resulting in better performances than hand-coded approaches. Nonetheless, careful design and engineering are required to define rewards to effectively and efficiently train RL agents. Many real-world tasks can be encoded by sparse rewards, which are in general easy to define but result challenging to optimize with RL. Instead, in robotics and other fields dense rewards are highly adopted in robotics to guide the exploration of autonomous agents [13] with the drawback of a more elaborate design. Indeed, dense reward functions can be difficult to tune, causing numerical instability, and misalignment w.r.t. the task objectives. Thus, automated or guided reward generation methods are interesting for both academic research and practical reasons. Several recent publications have tried to handle the design of dense reward functions in an automated fashion by using state-of-the-art coding LLMs, such as GPT-4 [14]. Their remarkable skills in code writing using natural language context allow non-expert users to define goals for RL agents as simple human-readable text. Such approaches reached remarkable performances in RL reward design in dexterous robotics manipulation [15, 16, 17, 18, 19], locomotion [20, 21], navigation [22], and autonomous driving [23]. Other applications of LLMs and VLMs for RL in robotics include approaches to guide the exploration process of RL agents [24], imitation [25, 26], online fine-tuning [27], and Preference-based RL (PbRL) [28]. Remarkably, [29] employs a VLM as the backbone of a Deep policy, and RL training is used to fine-tune the policy performance for various tasks in simulated environments. Current solutions for robotics manipulation, including [15, 7, 17, 18, 19], demonstrate few examples of applications in real-world settings, most requiring multiple training stages using feedback from human experts.
@@
In this work, we propose Autonomous Reinforcement learning for Complex Human-Informed Environments (ARCHIE). ARCHIE is a practical automatic RL pipeline for training autonomous agents for robotics manipulation tasks, in an unsupervised manner. ARCHIE employs GPT-4 —a popular pre-trained LLM— for reward generation from human prompts. We leverage natural language descriptions to generate reward functions via GPT-4, which are then used to train an RL agent in a simulated environment. Our approach introduces a formalization of the reward function that constrains the language model’s code generation, enhancing the feasibility of task learning at the first attempt. Unlike previous methods, we also utilize the language model to define the success criteria for each task, further automating the learning pipeline. Moreover, by properly formalizing the reward functions in shaping and terminal terms, we avoid the need for reward reflection [15] and multiple stages of training in RL [18]. This effectively reduces training time for RL agents, which is crucial, as training time impacts not only applicability of such methods, but also their energy and resource consumption. This results in a streamlined, one-shot process translating the user’s text descriptions into deployable skills. ARCHIE relies on the reasonable assumption that the manipulator’s kinematic and dynamic models are available and that RL agents are trained in a simulated environment. We demonstrate the effectiveness of the reward generation in simplified simulated environments and with a selection of real-world single-arm and bi-manual manipulation tasks on an ABB YuMi collaborative robot, shown in Figure 5. Policies trained in simulation are demonstrated on the real robot in laboratory settings 111videos available at https://youtu.be/kaTxZ4oqD6Y, thus avoiding the use of safe RL methods..
@@ II-A LLM-assisted RL
In [24], the authors propose the use of a VLM and an LLM to guide policy learning in the context of robotics manipulation. The VLM is used as an open-vocabulary object detector to get the observation space of the RL policy, while the action space is defined w.r.t. the reference frame of the objects. The LLM is used to program both a high-level and a low-level policy to generate affordance maps [3] according to the visual input. The LLM policies are used during the training of a Soft Actor-Critic (SAC) [30] agent with an $\varepsilon$-greedy exploration approach. In [25, 26], the authors use the LLMs to define expert agents to guide the RL training with imitation. The approach in [25] relies on the availability of pick/place primitives that make the RL agent’s action space discrete. The LLM policies generate expert demonstration data that bootstrap the training of an SDQfD [31] agent. In [26], a variation of Twin Delayed Deep Deterministic Policy Gradient (TD3) [32] is proposed, to accommodate an imitation loss in the actor’s loss. The imitation data is collected during the training process with a robot controller programmed with an LLM. In [27] the researchers used Cal-QL [33], a two-step policy training composed of (i) an offline stage where a policy is pre-trained to imitate various demonstrations, (ii) an online stage where the policy is fine-tuned to solve specific tasks while using a VLM to generate rewards directly from visual observations. In [28] authors incorporate PEBBLE [34] with natural language user feedback on the iteratively learned robot policies during the PbRL training. Preferences of multiple LLMs are collected to compute a scoring of the performance of agents.
@@ II-B LLM for reward generation
Most notably, Eureka [15] is one of the most recent contributions in reward generation for RL. It employs GPT-4 to generate dense reward functions in the form of Python code by using the environment’s source code and a natural-language task description as context. Then, it iterates the training and reward reflection by querying the LLM with feedback from the previous episode to improve the performance of the agents. Such a strategy has proven to be effective in simulation, where it can accomplish various robotic tasks that cannot be solved in a sparse reward setting. Remarkably, the reward functions generated by GPT-4 prove to be more robust than rewards generated by humans. Recently, Eureka was applied in a real-world robotic locomotion task [35], but to the best of our knowledge, not to robotic manipulation. Similarly, in [18] and [19], the authors use an LLM to design reward functions based on natural language descriptions of the task and environment. The initial reward function design is then used to train an RL agent, and the results are fed back to the LLM for self-refinement. This iterative process continues until the reward function meets the desired performance criteria. Notably, [19] presented results in real-world dexterous manipulation tasks. In [16], authors use Offline RL [36] to train language-conditioned robot behaviors from large annotated offline datasets that outperform language-conditioned imitation learning techniques. The method exhibits zero-shot generalization to unseen natural language commands in real-world settings. The work in [17] leverages LLMs and VLMs to extract knowledge about object semantics and common-sense task relevance to autonomously and continuously generate new skills through a self-guided learning cycle. In this cycle, the VLM-powered agent autonomously decides the learning task and prepares a natural language description followed by a cost or reward function. Remarkably, it also generates an appropriate simulated environment using a dictionary of available objects. Tasks are solved with RL, Model Predictive Control (MPC) [37] or other Optimal Control strategies. With this approach, the agent builds a plethora of skills for a variety of self-proposed tasks.
@@ III Notations
RL algorithms aim at optimizing the policy $\pi$ to maximizes the expected discounted sum of rewards $R_{0}=\mathbb{E_{\pi_{\phi}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})\right]$ of a specific MDP. The main mathematical tool designed to solve RL tasks is action-value function $Q$:
@@ IV-B Initialization: Natural Language to Rewards
Given a task for the robot, the LLM is used to write a new Python class that extends the base class by defining the dense reward function to guide the agent’s exploration towards the task goal. Unlike previous publications, we require the LLM to implement the termination conditions for the defined task, including a success condition that evaluates whether the task is solved and a failure condition that assesses whether the agent has failed. This is done to reduce the ambiguity in task definitions in natural language, as the geometrical context is very complex to infer from the environment’s code. Moreover, the failure condition reduces the amount of exploration required by the agents, since failing states are associated with low rewards and are not interesting to explore. In this way, the agent is encouraged to avoid failure instead. In Figure 1 (yellow block) we report an example of user task description, including clear geometrical definitions of success and failure for a cube pushing task. With clear geometrical indications, writing the termination conditions becomes a trivial programming task, which the LLM can solve.
@@
  • Introduction, in which the overall LLM’s code generation task is described, including the expected format of the output Python code;
@@
  • Coding Context, including the Python code and documentation of the base environment class, and the instructions to implement the dense reward formalization in Section IV-C. The documentation is an important part of the context as it explains how to use the APIs to register contacts and poses of the objects useful for defining the reward components.
  • RL Context: general indications on reward functions for manipulation tasks, including some examples.
The prompt design is overviewed in Figure 1 (blue box), and the full text is reported in the PDF appendix. As in previous work, the code generation instructions are generic and high level, as examples of RL code, including rewards, are present in modern LLM training data.
@@ IV-C Autonomous Skill Learning: Dense reward formalization
Let $r^{k}(s_{t},a_{t})$, $k=1,\dots,K$ be the $k$-th reward component of our dense reward function shaping, received after applying action $a_{t}$ in state $s_{t}$. These terms are generated by the LLM using the prompt engineering defined in Section IV-B and overviewed in Figure 1. Arbitrarily, the reward components $r^{k_{+}}(s_{t},a_{t})$, $k_{+}\in\boldsymbol{K_{+}}\subseteq\{1,..,K\}$ are positive and are called bonuses. Due to the available context, the shaping terms are generally coherent with the task requirements. Nevertheless, there is no guarantee that the different generated terms are numerically coherent with each other, as discussed in other works. In fact, the authors in [15, 18] propose iterative processes in which the LLM evaluates the results of the trained agents and adjusts the weights of the different addends. In an effort to avoid this iterative process, we introduce a clearly defined terminal reward term $R_{F}(s_{t},a_{t})$ and two binary classifiers $\Phi(s_{t})$ and $\Gamma(s_{t})$, which implement respectively the success and failure conditions checks on state $s_{t}$.
$r^{\prime}(s_{t},a_{t},s_{t+1})=\sum_{k=1}^{K}r^{k}(s_{t},a_{t})+R_{F}(s_{t},a_{t})\cdot\Phi(s_{t+1})$ (3)
@@
To ensure that the terminal reward is sufficiently high in any terminal state $s_{t^{\prime}+1}$, the sufficient condition is
$r^{k_{+}}(s_{t^{\prime}},a_{t^{\prime}})\geq r^{k_{+}}(s_{t},a_{t},)\quad\forall t=0,\dots,t^{\prime},{k_{+}}\in\boldsymbol{K_{+}}$ (5)
@@
The statement is true if $\sum_{k_{+}\in\boldsymbol{K_{+}}}r^{k}(s_{t^{\prime}},a_{t^{\prime}})\leq 1$, otherwise it is sufficient to apply eq. 5 to the definition in eq. 4:
@@
$\sum_{t=0}^{t^{\prime}}\sum_{k_{+}\in\boldsymbol{K_{+}}}r^{k_{+}}(s_{t},a_{t})$ $>\sum_{t=0}^{t^{\prime}}r(s_{t},a_{t})\quad\Box$ (7)
In this way, we guarantee that the terminal reward for completing the task is always at least one order of magnitude greater than the cumulative rewards that the agent can collect during the trajectory before termination. For $\gamma<1$, this encourages the agent to move toward the terminal states rather than stagnating in high-rewarding non-terminal states, which is likely to happen if the shaping components are not balanced. The formalization is enforced as a hard programming constraint in the LLM’s prompt, as reported in Figure 1 (blue box), and in the PDF appendix.
Remark: this hard constraint on the code generation is the main difference w.r.t. previous work, e.g. Eureka [15].
Figure 2: Cube push example trainings. SAC trained with rewards generated by GPT-4, with and without the proposed formalization.
@@ IV-D Autonomous Skill Learning: Policy Learning
The pseudocode of the policy learning algorithm is reported in the PDF appendix (Algorithm 1), it follows the notations used in [30, 32] and related research. Success and failure conditions $\Phi$ and $\Gamma$ are used to terminate episodes. In the simulated environment, the agent’s state is defined as a vector containing the pose and linear and angular velocities of the robot’s end-effectors, the position and velocity of the grippers’ joint, and the pose of the objects in the scene. The agent’s action is defined as the reference linear and angular velocities of the end-effectors and the reference velocity of the grippers’ joint. This definition of state and action spaces allows the agent to directly correlate actions with changes in the state and reward. For single-arm tasks, the state and action spaces are limited to drive the left arm, while the right arm is fixed to a constant joint configuration.
Figure 3: Top: Success rates for SAC agent training on the selected benchmark tasks. The proposed reward formalization is compared with the performance of the rewards completely defined by Eureka at step 0 with GPT-4. For each task, 10 rewards are generated and each reward is tested with 10 different random seeds. Plots show median and 25-75 percentiles. Bottom: The agents performing the tasks. In the 2D environment, the agent (ball) is purple when holding the object (rectangle), while green otherwise.
V Experiments
In this section, we present the experimental evaluations performed to validate the proposed method. First, we validate the effectiveness of the reward formalization described in Section IV-C on three tasks on a simple 2D and 3D simulation. Second, we present the results of the Autonomous RL pipeline for 10 different robotics manipulation tasks, including demonstrations on the real setup.
V-A Benchmark Environments
To assess the effectiveness of the proposed reward formalization, we implemented simplified simulations that mimic the simplified dynamics of a robotics manipulation scene. In a first environment, the agent controls the movements of a point on a 2D vertical plane, a rectangular object is present on the “floor” of the plane. When in proximity to the object, the agent can perform grasping, which constrains the object to move solidly with the agent, until it stops the grasping, releasing the object. We consider three simple manipulation tasks: (i) a grasp and lift task, (ii) a grasp and slide task, (iii) a placing task. Figure 3 (bottom) shows an example execution for each task in the 2d environment. In the first two tasks, the agent is initialized to a constant position $x_{0}$, while the object is initialized in a position $x_{r}\sim p(x_{r})$. In the third task, the object and the agent are initialized next to each other, randomly at the top of the environment, $(x_{0},x_{r})\sim p(x_{0},x_{r})$.
In a second environment, implemented with 3D bullet simulation engine, the agent controls the movements of a small box, and can interact with a bigger box, over a narrow eleveted plane, with pushing actions. In this environment we consider a pushing task with constraint: the agent is required to push the object from one side of the surface to the other, without pushing the object from the surface. Figure 3 (bottom) shows an example of task execution in the 3D simulation.
For each task, we train agents with ARCHIE and a baseline reward generation. The selected baseline is the first step of reward generation of Eureka [15] and similar works [18], which is equivalent to the reward generation of ARCHIE, minus the formalization defined in Section IV-C. This comparison serves to highlight the improved learning stability due to the reward formalization.
We generate 10 rewards for both reward generations, using the same prompt instructions, except for the part that defines the formalization of the reward. The prompt structure and the natural language task definitions are available in the PDF appendix. We train 10 SAC agents for each reward using different random seeds and monitor the success rates of the trained policies. We use a single training environment ($N=1$) per agent. The results are reported in Figure 3 (top), where we show the statistics of the success rate curves of the agents trained by ARCHIE and the statistics of the rewards generated by Eureka (step 0).
In the training horizons considered, the agents trained by ARCHIE with the proposed formalization consistently solve all tasks, while the agents trained with Eureka (step 0) are much less consistent. This suggests that the proposed formalization effectively stabilizes policy learning toward task completion.
Running example: We remark one example of reward generated by ARCHIE for the Cube Push task in the 3D environment (Figure 3 bottom far right):
$r(s_{t},a_{t},s_{t+1})=-d+10\cdot contact+x_{t}+\underbrace{R_{F}(s_{t},a_{t})\Phi(s_{t+1})}_{\text{terminal term}},$ (8)
Here $\Phi(s_{t})=\text{step}(x_{t}-table\_lenght)$, $d$ is the distance between the two cubes, $contact$ is an observed variable which is true when the cubes are in contact, and $x_{t}$ is the x coordinate of the blue cube. This is a typical example of reward with correct structure and unbalanced weights, since $d$ and $x_{t}$ are both in the order of the unit, while the weight of $contact$ is one order of magnitude greater. The Python code that implements this reward can be found in the PDF appendix. In Figure 2 we show the learning curves of SAC agents under the reward in Equation 10, as well as the same reward without the terminal term. Indeed, the agents trained with the terminal reward all converge to successful solutions, instead the other agents fail. Again, this further highlights the benefits of the proposed formulation.
V-B Robotic Manipulation
We select 10 robotic manipulation tasks, each identified by a codename, the detailed task descriptions used for the LLM’s prompts are included in the PDF appendix. Single arm tasks: (i) CubeSlide: push a blue cube close to a red cube; (ii) CubePick: pick the blue cube without hitting the red cube next to the first cube; (iii) CubeStack: the robot is holding a blue cube, stack the blue cube over the red cube; (iv) CubesTowerSlide: push two stacked cubes 10cm to the right; (v) VialGrasp: Grasp a vial from a vial holder; (vi) CentrifugeInsertion: the robot is holding the vial, insert the vial in a centrifuge. Dual arm tasks: (i) BimanualPenInTape: one gripper is holding a pen and the other is holding a tape, insert the pen in the tape’s hole; (ii) CubeInCup: one gripper is holding a cube and the other is holding a cup, place the cube into the cup; (iii) BimanualBoxLift: Lift a box with both grippers; (iv) BimanualHandover: one gripper is holding a box, the other is empty, hand over the box to the empty gripper. For each task, we define the natural language task description, as described in Section IV-B, which we use to generate a reward with ARCHIE’s proposed reward generation. The same descriptions are exploited to generate code policies [41] with GPT-4, using the same interface as the RL policies.
Figure 5: The YuMi robot performing some of the selected tasks.
Each reward function, including the baseline, is used to train 3 SAC agents, for each task, using different random seeds. Each agent explores $N=4$ environments in parallel, as described in Section IV-D.
The training results are reported in Figure 4, where we compare the agents in terms of success rate. The figure includes the average success rates of 10 different code policies for each task. The code policies do not solve any of the selected tasks, while the rewards generated by ARCHIE guide the agents toward task completion in all tasks, reaching high success rates in most tasks. In some tasks, the RL training under ARCHIE presents some instability, which could be due to the discontinuity in the reward function. Instead, the distance-based baseline rewards only solve a few of the tasks, with low success rates. These results suggest that the reasoning capabilities of the LLM are adequate to select the reward terms related to the objective task, while it would practically struggle to define the low-level skills required to solve these tasks without ad-hoc interfaces. Moreover, a proper mathematical structure in the formalization of the rewards is fundamental to exploit the code generation capabilities of the LLM.
Figure 5 depicts the YuMi robot performing some of the tasks in the lab settings, while the experiments are shown in the demonstration video. The robot is controlled with a joint position controller, tracking the trajectories performed by the policy trained in simulation. The gripper is operated with a combination of force/position control.
VI Discussion and Conclusions
In this work, we proposed ARCHIE, an unsupervised pipeline leveraging GPT-4 to automate reward function generation for Reinforcement Learning (RL) in robotic manipulation. By utilizing Large Language Models (LLMs), our method translates human-readable task descriptions into structured reward functions, enabling training of RL agents in simulated environments with minimal manual intervention. Unlike previous methods, our approach integrates both reward generation and task success criteria formulation, creating a one-shot process for autonomous skill acquisition, significantly reducing the required training time.
Furthermore, our proposed reward formalization ensures stable policy learning, at the first training attempt, by balancing shaping and terminal rewards, mitigating issues related to reward misalignment. In contrast, other works rely on reward reflection to iteratively improve the generated rewards, thus requiring several RL policy training attempts and prompts to the LLM to converge to a stable policy. In the benchmark environments, we showed that the proposed one-shot reward generation is much more effective than the first step of reward generation of Eureka [15], completely formalized by GPT-4, as also done in other works, such as [18].
While our approach significantly advances automated RL training, several avenues remain unexplored for future research. The introduction of reward function refinement through real-world feedback and human preferences may allow for the integration of the training pipeline in robot programming software. The integration of safe RL strategies in real-world fine-tuning could improve sim-to-real transfer for realistic applications. Finally, exploring more complex robotic scenarios and multi-agent learning setups could extend the applicability of our method to a broader range of real-world tasks. Our findings highlight the potential of LLMs in accelerating RL-based robotic learning, paving the way for more autonomous and efficient robot training paradigms.
Appendix
VI-A Pseudocode